# MSACL: Multi-Step Actor-Critic Learning with Lyapunov Certificates for Exponentially Stabilizing Control

### v2

@@ V-C Performance Evaluation and Stability Analysis
Evaluation results demonstrate MSACL’s superiority. Specifically, MSACL consistently achieves the highest AMCR across all six benchmarks. Regarding AMCC, MSACL attains the absolute minimum in the Pendulum, SingleCarTracking, and QuadrotorTracking tasks, while remaining competitive in others. Furthermore, MSACL maintains the highest RR across all radii. Notably, in the high-dimensional QuadrotorTracking task, MSACL is the only approach to achieve a 100% reach rate at the $0.01$ precision level, whereas baselines fail. These results demonstrate the speed and precision of MSACL in complex control systems.
@@
Since MSACL builds on the SAC framework, SAC serves as the primary baseline for our ablation study. As shown in Fig. 4 and Table II, the integration of the Lyapunov certificate significantly enhances both training performance and stability metrics compared to the standard SAC. This validates that embedding Lyapunov-based constraints into MERL effectively yields neural policies with superior stability guarantees for model-free control systems.
@@ V-D Evaluation of Robustness and Generalization
For stabilization tasks (VanderPol, Pendulum, DuctedFan, Two-link), we introduce process noise $\bm{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\sigma_{\text{env}}\cdot\mathbf{1}_{d})$ to the system dynamics, where $\mathbf{1}_{d}$ denotes the all-ones vector of dimension $d$, yielding the perturbed form:
@@
Robustness results for stabilization tasks in Fig. 6 show that all trajectories converge to the equilibrium $\mathbf{x}_{g}=\mathbf{0}$ despite parametric perturbations and process noise. Tracking task results (Fig. 7, 8) demonstrate that MSACL-trained policies effectively mitigate process noise within reasonable ranges. For SingleCarTracking, though trained on straight-line references, the policy generalizes well to circular and sinusoidal trajectories (Fig. 7c, 7d), with position error $e_{\mathrm{xy}}=\sqrt{e_{x}^{2}+e_{y}^{2}}$ remaining below 0.1 in steady state (Fig. 7e). For QuadrotorTracking, though trained on horizontal helical trajectories, the policy successfully tracks unseen vertical helix and Lissajous curves (Fig. 8c, 8d), with tracking error $e_{\mathrm{xyz}}=\sqrt{e_{x}^{2}+e_{y}^{2}+e_{z}^{2}}$ stabilizing around 0.02 (Fig. 8e), demonstrating superior generalization capability.
Figure 6: System trajectories under the MSACL-trained policies in stabilization tasks subjected to environmental parametric uncertainties and process noise: (a) VanderPol with state variables $\mathbf{x}=[x,\dot{x}]^{\top}$; (b) Pendulum with angular position $\theta$ and angular velocity $\dot{\theta}$; (c) DuctedFan with coordinates $x,y$ and orientation $\theta$; (d) Two-link with joint angles $q_{1}$ and $q_{2}$. All trajectories exhibit robust convergence to the neighborhood of the equilibrium $\mathbf{x}_{g}=\mathbf{0}$.
@@ V-E Sensitivity Analysis of the Multi-Step Horizon
This subsection examines the impact of multi-step horizon $n$ on MSACL performance across all benchmarks. Following the protocol in Section V-C, we evaluate policies trained with varying $n$ using performance metrics (AMCR, AMCC) and stability metrics (RR, ARS, AHS). For each benchmark, we conduct 100 experiments from fixed random initial states to compute these metrics. As summarized in Table IV, performance generally improves with larger $n$, as a longer horizon enables the policy to better identify the stability direction, i.e., the descent of $V_{\psi}$. This reduces training bias and reinforces stability guarantees.
@@ VI-A Discussion
This subsection elucidates the functional roles of MSACL components. As highlighted in Remark 4, MSACL augments the SAC framework with a stability-oriented critic, departing from standard RL that relies on exhaustive exploration to discover stabilizing behaviors. Instead, MSACL leverages a learned Lyapunov certificate to provide explicit, structured guidance for policy updates.
Theoretically, MSACL is grounded in Lemma 1, which establishes the equivalence between Lyapunov stability and exponential convergence. To operationalize this exponential stability requirement, the ESL mechanism (Definition 2) categorizes trajectory segments based on their contraction properties, transforming the stability condition into a supervised learning task. This ensures the synthesis of a high-fidelity Lyapunov certificate $V_{\psi}$ that accurately captures the underlying system dynamics. Such fidelity is crucial for providing the actor with theoretically sound, stability-aware gradients. Consequently, the stability advantage in (18) prioritizes actions that facilitate Lyapunov descent, effectively guiding the policy toward the region of stability.
While developed for stabilization and tracking, MSACL is extensible to safety-critical applications including autonomous vehicle trajectory planning [52], spacecraft proximity operations [53], and safety-guaranteed robotic manipulation [54, 55]. The integration of Lyapunov guarantees within MERL represents a step toward verifiable learning-based control.
@@ VI-B Conclusion
This article presents MSACL, a novel Lyapunov-based RL algorithm that integrates exponential stability into MERL and provides a framework for developing verifiable, learning-based policies in safety-critical applications. By embedding Lyapunov theory within the SAC architecture, MSACL achieves stable control for model-free systems while preserving maximum entropy exploration benefits. Comprehensive evaluation demonstrates consistent superiority of MSACL over established baselines in both performance and stability metrics, with remarkable robustness and generalization capabilities. Future work will explore extensions to partially observable systems, adaptive Lyapunov architectures, and cross-domain transfer learning for complex dynamical environments.

### v3

@@ V-C Performance Evaluation and Stability Analysis
Evaluation results demonstrate MSACL’s superiority. Specifically, MSACL consistently achieves the highest AMCR across all six benchmarks. Regarding AMCC, MSACL attains the minimum cost among all evaluated methods in the Pendulum, SingleCarTracking, and QuadrotorTracking tasks, while remaining competitive in others. Furthermore, MSACL maintains the highest RR across all radii. Notably, in the high-dimensional QuadrotorTracking task, MSACL is the only approach to achieve a 100% reach rate at the $0.01$ precision level, whereas baselines fail. These results demonstrate the speed and precision of MSACL in complex control systems.
@@
Since MSACL builds on the SAC framework, SAC serves as the primary baseline for our ablation study. As shown in Fig. 4 and Table II, the integration of the Lyapunov certificate enhances both training performance and stability metrics compared to the standard SAC. This validates that embedding Lyapunov-based constraints into MERL effectively yields neural policies with stability properties for model-free control systems.
@@ V-D Evaluation of Robustness and Generalization
For stabilizing tasks (VanderPol, Pendulum, DuctedFan, Two-link), we introduce process noise $\bm{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\sigma_{\text{env}}\cdot\mathbf{1}_{d})$ to the system dynamics, where $\mathbf{1}_{d}$ denotes the all-ones vector of dimension $d$, yielding the perturbed form:
@@
Robustness results for stabilizing tasks in Fig. 6 show that all trajectories converge to the equilibrium $\mathbf{x}_{g}=\mathbf{0}$ despite parametric perturbations and process noise. Tracking task results (Fig. 7, 8) demonstrate that MSACL-trained policies effectively mitigate process noise within reasonable ranges. For SingleCarTracking, though trained on straight-line references, the policy generalizes well to circular and sinusoidal trajectories (Fig. 7c, 7d), with position error $e_{\mathrm{xy}}=\sqrt{e_{x}^{2}+e_{y}^{2}}$ remaining below 0.1 in steady state (Fig. 7e). For QuadrotorTracking, though trained on horizontal helical trajectories, the policy successfully tracks unseen vertical helix and Lissajous curves (Fig. 8c, 8d), with tracking error $e_{\mathrm{xyz}}=\sqrt{e_{x}^{2}+e_{y}^{2}+e_{z}^{2}}$ stabilizing around 0.02 (Fig. 8e), demonstrating superior generalization capability.
Figure 6: System trajectories under the MSACL-trained policies in stabilizing tasks subjected to environmental parametric uncertainties and process noise: (a) VanderPol with state variables $\mathbf{x}=[x,\dot{x}]^{\top}$; (b) Pendulum with angular position $\theta$ and angular velocity $\dot{\theta}$; (c) DuctedFan with coordinates $x,y$ and orientation $\theta$; (d) Two-link with joint angles $q_{1}$ and $q_{2}$. All trajectories exhibit robust convergence to the neighborhood of the equilibrium $\mathbf{x}_{g}=\mathbf{0}$.
@@ V-E Sensitivity Analysis of the Multi-Step Horizon
This subsection examines the impact of multi-step horizon $n$ on MSACL performance across all benchmarks. Following the protocol in Section V-C, we evaluate policies trained with varying $n$ using performance metrics (AMCR, AMCC) and stability metrics (RR, ARS, AHS). For each benchmark, we conduct 100 experiments from fixed random initial states to compute these metrics. As summarized in Table IV, performance generally improves with larger $n$, as a longer horizon enables the policy to better identify the stability direction, i.e., the descent of $V_{\psi}$. This reduces training bias and reinforces stability properties.
@@ VI-A Discussion
This subsection elucidates the functional roles of MSACL components and discusses the fundamental limitations of our approach. Functionally, MSACL augments the SAC framework with a stability-oriented certificate, departing from standard RL that relies on exhaustive exploration to discover stabilizing behaviors. Instead, MSACL leverages a learned Lyapunov certificate to provide guidance for policy updates. By categorizing trajectory segments through proposed sampling mechanism, the exponential stability constraints are coupled with the training samples, resulting in an empirically accurate Lyapunov certificate. Consequently, the stability advantage explicitly prioritizes actions that facilitate Lyapunov descent, effectively guiding the policy toward the region of stability.
Despite its favorable empirical performance, particularly its strong effectiviness and high efficiency, it is crucial to acknowledge that MSACL, fundamentally rooted in model-free RL, lacks formal theoretical guarantees for system stability. In contrast, traditional frameworks such as cMDPs [9, 42] and Hamilton-Jacobi reachability analysis [10, 43] offer rigorous mathematical formulations that provide strict theoretical guarantees for system behavior. Nevertheless, these formal methods often struggle with the curse of dimensionality and high computational costs in unknown environments. Bridging the gap between the empirical scalability of deep RL methods like MSACL and the rigorous theoretical guarantees of formal mathematical frameworks remains an essential challenge for future research.
@@ VI-B Conclusion
We in this article presented MSACL, a multi-step Lyapunov-guided reinforcement learning algorithm that provides an efficient framework for stabilizing control tasks. By integrating a Lyapunov-oriented critic into an off-policy MERL architecture, MSACL enables rapid and robust stabilization for model-free systems while improving sample efficiency through data reuse and broader exploration. Extensive experiments across multiple benchmarks demonstrated the consistent empirical superiority of MSACL over established baselines in terms of performance, robustness, and generalization capabilities.
Despite these promising results, the learned Lyapunov certificates in MSACL remain empirical approximations and do not yet provide strict formal guarantees comparable to those of classical analytical methods. Therefore, an important direction for future work is to integrate MSACL with formal verification tools. In particular, we plan to combine the proposed multi-step RL framework with structured analytical Lyapunov certificate templates, with the goal of enhancing both theoretical rigor and practical reliability. By bridging the scalability and exploration efficiency of MSACL with the mathematical rigor of formal methods, we hope to develop learning-based control strategies that are not only highly efficient, but also equipped with provable stability and safety guarantees.